8g Ca++ contain what number of charges
1
Expert's answer
2012-07-10T07:41:45-0400
The molar mass of Calcium is 40 g/mol
Thus, 8 g of Ca correspond to n(ca) =
mCa)/M(Ca) = 8 g / 40 g/mol = 0.2 mol of Ca.
One Ca++ ion has 2 charges, so
0.2 mol of Ca++ contain 0.4 mol of charges.
One mol of charges contain 96485
coulombs, so 0.4 mol contain n(Ca++) * F = 0.4*96485 = 38594 C.
(F = Faraday
constant is the magnitude of electric charge per mole of electrons, F = 96485
C/mol)